OpenStudy (anonymous):

check my work?? 5.) kim is studying the sale of a particular brand of cereals from the year 2000 to 2010. She writes the following functions to model the sale of the cereal S(t) in million dollars after t years. S(t)=t^2+7t +69 Part a. what does the y-intercept of the graph of the function represent? the sale of the cereal at the start Part B: What is the reasonable domain of the graph of the function? The domain is the time from 2000 to 2010 Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? t^2 + 7t + 69 4^2 + 7(4) +69 16 + 28 + 69=113

OpenStudy (anonymous):

does it look right?@thomaster

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

B and C are wrong

OpenStudy (the_fizicx99):

A reasonable domain for this function would be: 200 ≥ x ≤ 2010 because she's studying from the year 2000 to the year 2010.

OpenStudy (anonymous):

what wrong with c?

OpenStudy (the_fizicx99):

To find the average rate of change apply change in y all over change in x. (y - y)/(x - x). In this case it's change in y all over change in t, (y - y)/(t - t)

OpenStudy (the_fizicx99):

Set it up, tag me when you're done simplifying, plug in t = 1 and t = 4 into your function

OpenStudy (anonymous):

S(t)=t^2+7t +69 S(1)=4^2+7(4) +69 s(1)=16+28+69 73? @tHe_FiZiCx99

OpenStudy (the_fizicx99):

How did you get 73?

OpenStudy (the_fizicx99):

16 + 28 + 69 = 113 So (4,113) And when you plug in 1 for t you get (1,77) so \(\ \sf \dfrac{113-77}{4-1} = \dfrac{36}{3} = 12\) So the average rate of change between the \(\ \sf 4^{th}\)year and \(\ \sf 1^{st}\)is 12.

OpenStudy (anonymous):

ohhhh okay I see can you help me with another one? @tHe_FiZiCx99

OpenStudy (the_fizicx99):

If I can, I'm getting off soon